OpenStudy (anonymous):

Car license plates consists of 7 digits. How many license plates can be made?

OpenStudy (shubhamsrg):

You have to choose digits out of 10. Can your combination of digits start with 0 ?

OpenStudy (anonymous):

No ,i think only from 1- 9.

OpenStudy (shubhamsrg):

0 is also a digit. And let us assume your combination does not start with 0. Then choose 7 digits from 0-9. From this, subtract the possible combinations starting with 0. Here we have also assumed digits do not repeat.

OpenStudy (anonymous):

Imagine a simple binary number. \[0000\]. Each digit can take the value of 1, or 0, meaning, you are limited to two numbers. How many combinations can we form with these two numbers (0 or 1), in 4 places? Well, it's \[2^4 = 2*2*2*2 = 16\]. 16 is the maximum number of combinations for a binary number of 4 places. You can take this concept and apply it to your problem, except this time, you have 9 numbers to choose from (1-9), and 7 places.

OpenStudy (anonymous):

Thank you guys so much, so it would be 8*8*8*8*8*8*8?

OpenStudy (anonymous):

Well, actually, if you have numbers 1-9, it would be \[9^7 = 9*9*9*9*9*9*9\]

OpenStudy (anonymous):

But as @shubhamsrg stated, 0 is a digit, so I imagine it will actually be \[10^7 = 10*10*10*10*10*10*10\]

OpenStudy (anonymous):

ahh alright i understand now! Thank you! :)